OpenStudy (haileyd):

Sales for a business were 3 million dollars more the second year than the first, and sales for the third year were double the sales for the second year. If sales for the third year were 38 million dollars, what were sales, in millions of dollars, for the first year?

OpenStudy (doodleduck8):

16 is If x = sales for the first year, then x + 3 = sales for the second year. Since sales for the third year were double the sales for the second year, sales for the third year = 2(x + 3). Sales for the third year were 38, so 2(x + 3) = 38. To solve this equation, you could first divide each side by 2 to get x + 3 = 19. Then, by subtracting 3 from both sides, x = 16.

OpenStudy (haileyd):

Well thanks. That looked complicated but it wasn't

OpenStudy (doodleduck8):

Lol, your're welcome

OpenStudy (doodleduck8):

Thanks for the medal!!

OpenStudy (haileyd):

Welcome